A large LED screen is being installed at the Hotel Denouement as requested by Frank Denouement, the manager of the family-owned hotel. His identical brothers, Ernest and Dewey, went and installed the heavy screen (dubbed the Denouement Screen) to their lobby using four (4) 0.5-cm round bolts. If the Denouement Screen weighed 42.8 kg, how much shear stress do the bolts experience?
1
Expert's answer
2020-06-22T09:28:44-0400
"\sigma=\frac{mg}{4(0.25\pi d^2 )}"

"\sigma=\frac{42.8\cdot 9.8}{\pi( 0.005^2 )}=5.34\cdot 10^6\frac{N}{m^2}"